OpenStudy (anonymous):

Factor each trinomial below. Please show your work and check your answer. x2 – 8x + 15 a2 – a – 20 a2 – 5a – 20 a2 + 12ab + 27b2 2a2 + 30a + 100 Pleasepleaseplease

OpenStudy (anonymous):

is it x^2 or 2x

OpenStudy (anonymous):

(-5+x) (-3+x), (-5+a) (4+a), -20-5 a+a^2, (a+3 b) (a+9 b), 2 (5+a) (10+a)

OpenStudy (anonymous):

x2 – 8x + 15 = (x-5)(x-3) a2 – a – 20 = (a-5)(a+4) a2 – 5a – 20 = not fractor a2 + 12ab + 27b2 = (a+9b)(a+3b) 2a2 + 30a + 100 = (2a+10)(a+10)